Grade 11Trigonometry

Prove that Cos(b-c)+Cos(c-a)+Cos(a-b)=-3/2 if Cosa+Cosb+Cosc=0 and Sina+Sinb+Sinc=0

Vinay R
15 Years agoGrade 11

1 Answer

jitender lakhanpal
15 Years ago

Dear vinay,

 

squaring the equations 

Cosa+Cosb+Cosc=0    and  

Sina+Sinb+Sinc=0

 we get 

      cos2 a + cos2 b+ cos2 c + 2cosacosb+2cosbcosc+2cosacosc = 0

and similarly sin2 a +sin2  b +sin2  c+ 2 sinasinb+2 sinasinb+2 sinasinb=0

adding the above two equations we get 

3 + 2(sinasinb+cosacosb+sinasinb+cosacosb+sinasinb+cosacosb)=0

and by formula sinasinb + cosacosb = cos(a-b) and replacing

we get   

 

 Cos(b-c)+Cos(c-a)+Cos(a-b)=  -3/2  hence proved

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar: Click here to download the toolbar.. 

jitender

askiitians expert